Swifty Company sells 8% bonds having a maturity value of $3,000,000 for $2,772,550. The bonds are dated January 1, 2020, and mature January 1, 2025. Interest is payable annually on January 1. Determine the effective-interest rate. The effective-interest rate
Answer #1
Calculation of discount
Bond face value or maturity value 3000000
Less: Bond discount/issue value 2772550
Discount 227450

Bond period  January 1, 2020 to January 1, 2025 so,

No.of year 5

Discount per year=Discount/No. of year

=227450/5

=45490

Discount per year 45490
interest@8% 240000
Total 285490

average of discount price and face value=(face value+discounted value)/2

=(3000000+2772550)/2

average of discount price and face value= 4386275

Effective interest rate =285490/4386275*100=6.51%